# Learning to See and Act: Task-Aware Virtual View Exploration for Robotic Manipulation

### v4

@@ 4.1 Datasets and Experimental Setups
Simulation. Our experiments are conducted on RLBench dataset [james2020rlbench] and RLBench-OG.
RLBench is a multi-task learning benchmark implemented in the CoppeliaSim simulator featuring a 7-degree-of-freedom (7-DoF) Franka Emika Panda robot in a tabletop environment. Its observations consist of RGB-D images ($128\times 128$) captured from four fixed views, including frontal, left shoulder, right shoulder, and wrist-mounted. The action space represents the end-effector pose as $\mathbf{a}=(\mathbf{p}_{\text{xyz}},\mathbf{q}_{\text{rot}},g_{\text{gripper}},c_{\text{collision}})\in\mathbb{R}^{7}\times\{0,1\}^{2}$, where $\mathbf{p}_{\text{xyz}}\in\mathbb{R}^{3}$ denotes Cartesian coordinates, $\mathbf{q}_{\text{rot}}\in\mathbb{R}^{4}$ is the quaternion rotation, $g_{\text{gripper}}\in\{0,1\}$ controls gripper state, and $c_{\text{collision}}\in\{0,1\}$ indicates collision. Following previous works [goyal2024rvt2, ze2023gnfactor], we evaluate our method under two standard configurations: (1) a multi-view setup consisting of 18 manipulation tasks (2–60 variations each) with all four camera views, and (2) a single-view setup including 10 tasks observed only from the front RGB-D camera. The specific tasks are provided in the supplementary materials.
@@
Diffusion Policy [chi2023diffusion] $23.8\pm 1.9$ 3.0 $27.4\pm 0.8$ $3.4\pm 1.1 22$$9\pm 0.8 22.$$5\pm 0.7 22.$$2.6\pm 0.5$ $24.4\pm 1.4$ $4.9\pm 1.2 22$$2.2\pm 1.2$
ARP [zhang2025autoregressive] $63.7\pm 6.1$ 1.8 $73.0\pm 1.6$ $52.6\pm 1.0$ $59.8\pm 0.8$ $62.7\pm 0.5$ $61.3\pm 1.0$ $62.4\pm 1.4$ $68.1\pm 1.6$ $69.7\pm 1.4$
TVVE(Ours) $67.0\pm 6.2$ 1.1 $75.0\pm 1.2$ $58.0\pm 1.1$ $63.7\pm 1.1$ $4.6\pm 0.7 66$$6.8\pm 0.9$ $60.2\pm 0.6$ $74.3\pm 0.9$ $73.2\pm 0.9$
@@ 4.2 Comparison with State-of-the-art Methods
To comprehensively evaluate the effectiveness of our  TVVE, we conduct comparisons with a broad range of state-of-the-art approaches, including Diffusion Policy [chi2023diffusion], C2F-ARM-BC [james2022coarse], PerAct [shridhar2023perceiver], Hiveformer [guhur2023instruction], GNFactor [ze2023gnfactor], RVT [goyal2023rvt], RVT2 [goyal2024rvt2], Act3D [gervet2023act3d], 3D Diffuser Actor [ke20253d], and ARP [zhang2025autoregressive]. We evaluate them both quantitatively and qualitatively across three benchmarks: RLBench, RLBench-OG, and real-world manipulation tasks.
Comparisons on RLBench. Tables 1 and 2 present the quantitative results on RLBench [james2020rlbench] under the multi-view and single-view setups, respectively. Our proposed TVVE consistently achieves significant improvements in average success rate across both settings. Specifically, TVVE attains an average success rate of 86.6% under the multi-view setup, surpassing the previous state-of-the-art method ARP [zhang2025autoregressive] by approximately 5%. Notably, substantial gains are observed in tasks such as Insert Peg (from $65.6\%$ to $98.0\%$) and Sort Shape (from $36.0\%$ to $62.0\%$). A qualitative example of Insert Peg is illustrated in Fig. 4(a)(b). It shows that compared to ARP [zhang2025autoregressive], which relies on fixed viewpoints, our TVVE dynamically renders more informative perspectives, effectively mitigating occlusions and collisions while providing richer context for more accurate action prediction, validating that dynamic “seeing” underpins robust “acting.” Moreover, as indicated by the lower average rank of 2.17 in Table 1, TVVE demonstrates superior overall performance across all 18 tasks, highlighting its enhanced generalization capability, which is primarily attributed to the proposed TaskMoE design.
@@
Comparisons on Real-World Tasks. In Dobot robot experiment, compared to the baseline method Diffusion Policy (DP) [chi2023diffusion], TVVE achieves superior performance across all tasks, attaining a significantly higher average success rate of 88.0% versus 68.0% for DP. Notable gains are observed in tasks such as “Push Buttons” (+30%) and “Stack Bowls” (+20%), underscoring the effectiveness of integrating task-aware visual planning into the policy framework. A qualitative example of Pick Grape is illustrated in Fig. 4(c). In Franka robot experiment, TVVE achieves the highest average success rate of 78.0%, outperforming both Diffusion Policy (58.0%) and ARP (72.0%). These results further demonstrate its strong adaptability in real-world scenarios. We report the success rates based on 10 trials for each task, with detailed results provided in Table 4. More real-world results are shown in the supplementary material.
@@
Effectiveness of MVEP. The Multi-Viewpoint Exploration Policy (MVEP) is designed to dynamically explore optimal camera viewpoints that reduce occlusion and provide more informative visual feature for improved robotic manipulation. To evaluate the effectiveness of MVEP, we replace the explored viewpoints with either fixed viewpoints used in RVT2 [goyal2024rvt2] or randomly selected ones. As shown in Table 5, the manipulation success rate drops from $86.6\%$ to $83.3\%$ when using fixed viewpoints, and further plummets to $8.9\%$ when using random viewpoints. This significant degradation demonstrates that the learned viewpoint exploration policy in MVEP effectively identifies informative perspectives, which are crucial for accurate perception and robust action execution.
@@ 5 Conclusion
Our TVVE framework establishes that dynamic view planning and task-aware representation learning significantly advance robotic manipulation capabilities. The MVEP module’s viewpoint optimization effectively overcomes occlusion limitations in fixed-viewpoint systems, while TaskMoE’s specialized feature extraction mitigates multi-task interference. These contributions enhance performance across diverse manipulation challenges and enable meaningful generalization to novel tasks.
However, the framework has some limitations: multi-view re-rendering slightly increases inference latency, and reliance on accurate global point clouds makes it challenging to handle reflective or transparent objects. Future work could integrate multi-sensor data and domain adaptation to improve real-world robustness.
@@
  • Appendix D: Implementation details and information about the RLBench [james2020rlbench] and RLBench-OG. Additionally, we provide more visualized results, with corresponding demo videos included in the supplementary materials.
@@ Appendix B. More design details of TVVE
To handle heterogeneous manipulation tasks, we introduce a task-aware Mixture-of-Experts (TaskMoE) that routes visual tokens through (i) a task–instruction–aware gate layer and (ii) a per-gate expert selector. Unlike prior MoE work [shazeer2017outrageously, wangsparse], TaskMoE decouples tasks from gates, decouples gates from experts, and adds semantic/entropy regularizers to avoid collapse.
@@
In the RLBench dataset experiments, we select RGB-D images from four fixed viewpoints: front, right_shoulder, left_shoulder, and wrist as inputs. Following the approach of RVT-2 [goyal2024rvt2], we convert the depth maps from all viewpoints into point clouds in a global coordinate system using the intrinsic and extrinsic camera parameters. Subsequently, we perform multi-view point cloud aggregation and remove redundant points outside the workspace to obtain the final scene point cloud, as illustrated in Fig. 8.
@@ Appendix C. More Experimental Details and Resutls
We compare with eleven state-of-the-art baselines on RLBench, RLbench-OG and Real-world, including both 2D and 3D approaches. For 2D methods, we include Diffusion Policy [chi2023diffusion], which formulates robot policies as conditional denoising diffusion processes. The 3D baselines comprise voxel-based methods: C2F-ARM-BC [james2022coarse], which predicts actions in a coarse-to-fine manner via Q-value estimation; PerAct [shridhar2023perceiver], which detects actions through global self-attention; Hiveformer [guhur2023instruction], which employs attention across historical features; and GNFactor [ze2023gnfactor], which co-optimizes a neural scene representation with a PerAct-based policy. We also consider point-based methods: PolarNet [chen2023polarnet], which computes dense point representations, and multi-view approaches: RVT [goyal2023rvt], which fuses multi-view predictions via 3D back-projection, and its successor RVT2 [goyal2024rvt2], which enhances precision and efficiency. Furthermore, we include Act3D [gervet2023act3d], which uses coarse-to-fine 3D featurization (retrained by us for a fair comparison), 3D Diffuser Actor [ke20253d], a 3D diffusion-based policy, and ARP [zhang2025autoregressive], an autoregressive policy generating hybrid action sequences via a Chunking Causal Transformer.
@@
In the RLBench Multi-view setup experiments, during the supervised pre-training phase, our experimental hyperparameter settings comply with APR[zhang2025autoregressive]. In the offline reinforcement learning phase for MVEP, our hyperparameter configurations are detailed in Table 10.
@@
We conducted an evaluation of TVVE, ARP, and Diffusion Policy on the RLBench-OG benchmark. Detailed performance across 10 tasks under 8 variations is presented in Tables 13, 14, and 15. The reported success rates correspond to the mean $\pm$ standard deviation from three independent trials.
@@ Appendix D. Implementation details and information about the RLBench and RLBench-OG
The Generalization Suite evaluates robustness to environment-conditioned variations. Based on the same ten tasks, we construct six types of environment variations, each modifying exactly one factor while keeping all others unchanged. Following the pipeline from the COLOSSEUM [pumacay2024colosseum], we specify variation types using yaml configuration files and data collection procedures via json metadata.

### v5

@@ 4.1 Datasets and Experimental Setups
Simulation. Our experiments are conducted on RLBench dataset [22] and RLBench-OG.
RLBench is a multi-task learning benchmark implemented in the CoppeliaSim simulator featuring a 7-degree-of-freedom (7-DoF) Franka Emika Panda robot in a tabletop environment. Its observations consist of RGB-D images ($128\times 128$) captured from four fixed views, including frontal, left shoulder, right shoulder, and wrist-mounted. The action space represents the end-effector pose as $\mathbf{a}=(\mathbf{p}_{\text{xyz}},\mathbf{q}_{\text{rot}},g_{\text{gripper}},c_{\text{collision}})\in\mathbb{R}^{7}\times\{0,1\}^{2}$, where $\mathbf{p}_{\text{xyz}}\in\mathbb{R}^{3}$ denotes Cartesian coordinates, $\mathbf{q}_{\text{rot}}\in\mathbb{R}^{4}$ is the quaternion rotation, $g_{\text{gripper}}\in\{0,1\}$ controls gripper state, and $c_{\text{collision}}\in\{0,1\}$ indicates collision. Following previous works [15, 52], we evaluate our method under two standard configurations: (1) a multi-view setup consisting of 18 manipulation tasks (2–60 variations each) with all four camera views, and (2) a single-view setup including 10 tasks observed only from the front RGB-D camera. The specific tasks are provided in the supplementary materials.
@@
Diffusion Policy [9] (IJRR) $23.8\pm 1.9$ 4.0 $27.4\pm 0.8$ $3.4\pm 1.1 22$$9\pm 0.8 22.$$5\pm 0.7 22.$$2.6\pm 0.5$ $24.4\pm 1.4$ $4.9\pm 1.2 22$$2.2\pm 1.2$
ARP [55] (RA-L) $63.7\pm 6.1$ 2.5 $73.0\pm 1.6$ $52.6\pm 1.0$ $59.8\pm 0.8$ $62.7\pm 0.5$ $61.3\pm 1.0$ $62.4\pm 1.4$ $68.1\pm 1.6$ $69.7\pm 1.4$
RVT2 [15] (RSS) $64.5\pm 8.3$ 2.1 $72.8\pm 2.0$ $46.9\pm 0.4$ $60.8\pm 1.2$ $61.8\pm 1.1$ $64.0\pm 2.0$ $\mathbf{63.4\pm 0.2}$ $72.6\pm 0.6$ $\mathbf{74.0\pm 0.8}$
TVVE (Ours) $\mathbf{67.0\pm 6.2}$ 1.4 $\mathbf{75.0\pm 1.2}$ $\mathbf{58.0\pm 1.1}$ $\mathbf{63.7\pm 1.1}$ $\mathbf{64.6\pm 0.7}$ $\mathbf{66.8\pm 0.9}$ $60.2\pm 0.6$ $\mathbf{74.3\pm 0.9}$ $73.2\pm 0.9$
@@ 4.2 Comparison with State-of-the-art Methods
To comprehensively evaluate the effectiveness of our  TVVE, we conduct comparisons with a broad range of state-of-the-art approaches, including Diffusion Policy [9], C2F-ARM-BC [23], PerAct [45], Hiveformer [17], GNFactor [52], RVT [16], RVT2 [15], Act3D [14], 3D Diffuser Actor [25], and ARP [55]. We evaluate them both quantitatively and qualitatively across three benchmarks: RLBench, RLBench-OG, and real-world manipulation tasks.
Comparisons on RLBench. Tables 1 and 2 present the quantitative results on RLBench [22] under the multi-view and single-view setups, respectively. Our proposed TVVE consistently achieves significant improvements in average success rate across both settings. Specifically, TVVE attains an average success rate of 86.6% under the multi-view setup, surpassing the previous state-of-the-art method ARP [55] by approximately 5%. Notably, substantial gains are observed in tasks such as Insert Peg (from $65.6\%$ to $98.0\%$) and Sort Shape (from $36.0\%$ to $62.0\%$). A qualitative example of Insert Peg is illustrated in Fig. 4(a)(b). It shows that compared to ARP [55], which relies on fixed viewpoints, our TVVE dynamically renders more informative perspectives, effectively mitigating occlusions and collisions while providing richer context for more accurate action prediction, validating that dynamic “seeing” underpins robust “acting.” Moreover, as indicated by the lower average rank of 2.17 in Table 1, TVVE demonstrates superior overall performance across all 18 tasks, highlighting its enhanced generalization capability, which is primarily attributed to the proposed TaskMoE design.
@@
Comparisons on Real-World Tasks. In Dobot robot experiment, compared to the baseline method Diffusion Policy (DP) [9], TVVE achieves superior performance across all tasks, attaining a significantly higher average success rate of 88.0% versus 68.0% for DP. Notable gains are observed in tasks such as “Push Buttons” (+30%) and “Stack Bowls” (+20%), underscoring the effectiveness of integrating task-aware visual planning into the policy framework. A qualitative example of Pick Grape is illustrated in Fig. 4(c). In Franka robot experiment, TVVE achieves the highest average success rate of 78.0%, outperforming both Diffusion Policy (58.0%) and ARP (72.0%). These results further demonstrate its strong adaptability in real-world scenarios. We report the success rates based on 10 trials for each task, with detailed results provided in Table 4. More real-world results are shown in the supplementary material.
@@
Effectiveness of MVEP. The Multi-Viewpoint Exploration Policy (MVEP) is designed to dynamically explore better camera viewpoints that reduce occlusion and provide more informative visual feature for improved robotic manipulation. To evaluate the effectiveness of MVEP, we replace the explored viewpoints with either fixed viewpoints used in RVT2 [15] or randomly selected ones. As shown in Table 5, the manipulation success rate drops from $86.6\%$ to $83.3\%$ when using fixed viewpoints, and further plummets to $8.9\%$ when using random viewpoints. This significant degradation demonstrates that the learned viewpoint exploration policy in MVEP effectively identifies informative perspectives, which are crucial for accurate perception and robust action execution.
@@ 5 Conclusion
Our TVVE framework establishes that dynamic view planning and task-aware representation learning significantly advance robotic manipulation capabilities. The MVEP’s viewpoint optimization effectively overcomes occlusion limitations in fixed-viewpoint systems, while TaskMoE’s specialized feature extraction mitigates multi-task interference. These contributions enhance performance across diverse manipulation challenges and enable meaningful generalization to novel tasks. However, the framework has some limitations: multi-view re-rendering slightly increases inference latency, and reliance on accurate global point clouds makes it challenging to handle reflective or transparent objects. Future work could integrate multi-sensor data and domain adaptation to improve real-world robustness.
Acknowledgments
This work is supported by National Key Research and Development Program of China (2024YFE0203100), in part by the National Natural Science Foundation of China under Grant No.62572498 and No.62536010, in part by the open research fund of Pengcheng Laboratory under Grant 2025KF1B0050, in part by the Guangdong Basic and Applied Basic Research Foundation under Grant NO. 2025A1515011874. We also thank the National Supercomputer Center in Guangzhou for computational support.
@@ Acknowledgments
  • Appendix D: Implementation details and information about the RLBench [22] and RLBench-OG. Additionally, we provide more visualized results, with corresponding demo videos included in the supplementary materials.
@@ Appendix B. More design details of TVVE
To handle heterogeneous manipulation tasks, we introduce a task-aware Mixture-of-Experts (TaskMoE) that routes visual tokens through (i) a task–instruction–aware gate layer and (ii) a per-gate expert selector. Unlike prior MoE work [43, 50], TaskMoE decouples tasks from gates, decouples gates from experts, and adds semantic/entropy regularizers to avoid collapse.
@@
In the RLBench dataset experiments, we select RGB-D images from four fixed viewpoints: front, right_shoulder, left_shoulder, and wrist as inputs. Following the approach of RVT-2 [15], we convert the depth maps from all viewpoints into point clouds in a global coordinate system using the intrinsic and extrinsic camera parameters. Subsequently, we perform multi-view point cloud aggregation and remove redundant points outside the workspace to obtain the final scene point cloud, as illustrated in Fig. 8.
@@ Appendix C. More Experimental Details and Resutls
We compare with eleven state-of-the-art baselines on RLBench, RLbench-OG and Real-world, including both 2D and 3D approaches. For 2D methods, we include Diffusion Policy [9], which formulates robot policies as conditional denoising diffusion processes. The 3D baselines comprise voxel-based methods: C2F-ARM-BC [23], which predicts actions in a coarse-to-fine manner via Q-value estimation; PerAct [45], which detects actions through global self-attention; Hiveformer [17], which employs attention across historical features; and GNFactor [52], which co-optimizes a neural scene representation with a PerAct-based policy. We also consider point-based methods: PolarNet [8], which computes dense point representations, and multi-view approaches: RVT [16], which fuses multi-view predictions via 3D back-projection, and its successor RVT2 [15], which enhances precision and efficiency. Furthermore, we include Act3D [14], which uses coarse-to-fine 3D featurization (retrained by us for a fair comparison), 3D Diffuser Actor [25], a 3D diffusion-based policy, and ARP [55], an autoregressive policy generating hybrid action sequences via a Chunking Causal Transformer.
@@
In the RLBench Multi-view setup experiments, during the supervised pre-training phase, our experimental hyperparameter settings comply with APR[55]. In the offline reinforcement learning phase for MVEP, our hyperparameter configurations are detailed in Table 10.
@@
We conducted an evaluation of TVVE, RVT2, ARP, and Diffusion Policy on the RLBench-OG benchmark. Detailed performance across 10 tasks under 8 variations is presented in Tables 13, 14, 15, and 16. The reported success rates correspond to the mean $\pm$ standard deviation from three independent trials.
@@
basketball_in_hoop $\pm 0.0 100.0±0.010$$\pm 0.0$ $88.8\pm 1.6$ $6.0\pm 0.0 99$$9.2\pm 1.6$ $88.0\pm 0.0$ $\pm 0.0 100.0±0.010$$\pm 0.0$ $96.5\pm 4.8$
scoop_with_spatula $82.7\pm 7.5$ $84.0\pm 6.5$ $75.4\pm 2.6$ $4.0\pm 5.7 88$$8.0\pm 5.7$ $76.0\pm 4.0$ $98.0\pm 2.0$ $96.0\pm 0.0$ $85.5\pm 7.7$
take_plate_off_colored_dish_rack $98.7\pm 1.9$ $65.3\pm 6.8$ $60.0\pm 4.6$ $88.8\pm 3.9$ $84.8\pm 4.7$ $\pm 0.0 100.0±0.010$$\pm 0.0$ $100.0\pm 0.0$ $87.2\pm 15.2$
water_plants $72.0\pm 8.6$ $20.0\pm 5.7$ $12.0\pm 5.1$ $9.6\pm 9.3$ $3.0\pm 5.9 22$$2.0\pm 2.0$ $26.0\pm 2.0$ $32.0\pm 0.0$ $27.1\pm 18.3$
block_pyramid $4.0\pm 3.3$ $\pm 0.0 0.0±0.0$$\pm 0.0$ $\pm 0.0 0.0±0.0$$\pm 0.0$ $\pm 0.0 0.0±0.0$$\pm 0.0$ $0.0\pm 0.0$ $0.5\pm 1.3$
solve_puzzle $\pm 0.0 0.0±0.0$$\pm 0.0$ $9.3\pm 3.0$ $8.0\pm 3.6$ $12.0\pm 5.1$ $20.0\pm 0.0$ $16.0\pm 3.3$ $25.3\pm 3.8$ $11.3\pm 8.4$
take_usb_out_of_computer $\pm 0.0 100.0±0.010$$\pm 0.0$ $\pm 0.0 100.0±0.010$$\pm 0.0$ $95.2\pm 1.6$ $98.0\pm 2.0$ $98.7\pm 1.9$ $100.0\pm 0.0$ $99.0\pm 1.6$
close_drawer $\pm 0.0 100.0±0.010$$\pm 0.0$ $99.3\pm 1.5$ $97.6\pm 2.0$ $98.4\pm 2.0$ $96.0\pm 0.0$ $\pm 0.0 100.0±0.010$$\pm 0.0$ $98.9\pm 1.4$
straighten_rope $70.7\pm 10.0$ $0.0\pm 0.0$ $56.7\pm 5.8$ $37.6\pm 5.4$ $44.0\pm 5.1$ $36.0\pm 0.0$ $6.7\pm 5.0 88$$8.0\pm 4.0$ $52.5\pm 27.6$
toilet_seat_down $100.0\pm 0.0$ $0.0\pm 0.0$ $\pm 0.0 100.0±0.010$$\pm 0.0$ $100.0\pm 0.0$ $98.0\pm 2.0$ $\pm 0.0 100.0±0.010$$\pm 0.0$ $87.2\pm 33.0$
Task Mean $72.8\pm 2.0$ $46.9\pm 0.4$ $60.8\pm 1.2$ $61.8\pm 1.1$ $64.0\pm 2.0$ $63.4\pm 0.2$ $72.6\pm 0.6$ $74.0\pm 0.8$ $64.5\pm 8.3$
Table 14: Success Rates of RVT2 under Different Perturbations of RLBench-OG.
@@ Appendix D. Implementation details and information about the RLBench and RLBench-OG
The Generalization Suite evaluates robustness to environment-conditioned variations. Based on the same ten tasks, we construct six types of environment variations, each modifying exactly one factor while keeping all others unchanged. Following the pipeline from the COLOSSEUM [40], we specify variation types using yaml configuration files and data collection procedures via json metadata.